OpenStudy (anonymous):

Write the equation of a line in slope-intercept form passing through (4, 5) and (6, 8) A. y equals two-thirds x plus seven-thirds B. y equals three-halves x plus 11 C. y equals three-halves x minus 1 D. y equals three-halves minus 8

OpenStudy (tkhunny):

Have you considered actually writing the equation of the given line? You have two points.

OpenStudy (anonymous):

BY using y-y1=[(y2-y1)/(x2-x1)](x-x1) or y-y1=m(x-x1) m=(y2-y1)/(x2-x1)

OpenStudy (anonymous):

c?

OpenStudy (anonymous):

@chetan552

OpenStudy (tkhunny):

It appears that you are just guessing. Please show your work. No guessing will be required.

OpenStudy (anonymous):

Yes c is correct option @micahcarr76

OpenStudy (anonymous):

m=(8-5)/(6-4)=3/2 y-5=(3/2)(y-4) then y=(3/2)x-1

OpenStudy (anonymous):

Is your doubt clear @micahcarr76

OpenStudy (anonymous):

Don't forget to click on the button called "Best Answer" to give a medal to the person who helped you :)